-- Main.Zheng - 11 Dec 2012

6:45pm, Acqusition UDS mask.

7:10pm, Telescope has problem of tracking, and get fixed in 20min.

7:45pm, Start Science exposure, seeing ~1.5''

8:15pm, seeing ~1.4''

8:30pm, the instrument temperature going up.

8:45pm, changing to Wenwen's target, seeing ~1.7-2.0''

8:50pm, instrument temperature problem (too high), doing some checking before Wenwen's target get acquisition.

9:10pm, start to acquisition Wenwen's target J2327.

9:45pm, we change target to J0745.

9:55pm, we start the acquisition of J0745 in J band.

10:10pm, start the science script on J0745 in J band, seeing ~ 1.6''

10:50pm, start the science script on J0745 in K band, seeing ~ 1.4-1.5''

11:30pm, telescope tracking problem again ...

11:45pm, we re-acquisition to J0745 in K band, seeing ~1.9''

11:50pm, we start exposure on J0745 in K band.

12:30am, we change target to J0825 in J band, seeing < 1.3''

1:05am, we acquisition Greg's target. seeing ~1.0''

1:20am, we start the science exposure on Greg's target.

1:50am, we got 4x 300sec on Greg's target Abell 773, and moving to z7p7 Linhua's target. Seeing is 0.8-1.0''

2:00am, we start acquisition Linhua's z7p7 target.

2:10am, right after we done with the acquisition, the whole building's power off.

2:40am, we got electricity, but telescope get stuck at 75 degrees.

5:50am, we are waiting to do calibrations, and there are no chance to take science target after the building's power off.
Topic revision: r4 - 11 Dec 2012, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback